# PathoSyn: Imaging-Pathology MRI Synthesis via Disentangled Deviation Diffusion

### v1

@@ 1 Introduction
Pathological image generation is increasingly vital for medical image analysis, supporting data augmentation [frid2018synthetic, kebaili2023deep, koetzier2024generating], disease progression modeling [yoon2023sadm, puglisi2024enhancing], and clinical decision support [pinto2023artificial, comaniciu2016shaping]. However, acquiring balanced datasets is hindered by limited patient availability, ethical constraints, and high annotation costs. Generative models mitigate this scarcity by synthesizing realistic pathological images that complement real-world data. Beyond algorithmic utility, high-fidelity synthetic pathology aids in clinical intervention planning by illustrating plausible disease trajectories, surfacing rare edge cases, and enabling controlled evaluation of models prior to deployment [kim2022diffusion, shang2022learning, van2024synthetic, yoon2023sadm, puglisi2024enhancing]. Enhancing the fidelity and consistency of these images is therefore directly linked to developing safer, more reliable tools for clinical decision-making.
Despite growing interest, generating pathological images remains challenging due to the intrinsic asymmetry between anatomy and disease. The anatomical structure, including organ geometry and spatial layout, is largely stable for a given subject, whereas the pathological appearance is the main source of uncertainty, varying in intensity, texture, and clinical stage. The most clinically relevant variability is concentrated in the way disease perturbs a stable anatomical substrate. Existing generative models [koetzier2024generating, kazerouni2022diffusion] rarely make this distinction explicit. When operating directly in image space, anatomy and pathology are treated equally stochastic; the model must learn a high-dimensional distribution over all pixels simultaneously. This unnecessarily enlarges the solution space and permits the generator to introduce anatomical modifications in order to fit the data, thereby producing systematic distortions of anatomically stable structures. Consequently, current approaches fall into two suboptimal extremes: either they model the image as a holistic entity with no explicit separation [pinaya2022brain], or they enforce complete separation via masking, which identifies location, but overlooks internal lesion appearance and continuity with surrounding tissue.
@@
Although deep learning-based methods have significantly advanced visual fidelity compared to traditional rule-based algorithms, they have remained largely confined to the same image-space paradigm [jog2017random, ye2013modality]. Early applications of generative adversarial networks (GANs) synthesized lesions by modifying entire slices or volumes, sometimes conditioned on labels or coarse masks [goodfellow2020generative, isola2017image, wolterink2017deep, frid2018synthetic, yi2019generative]; these models can create sharp and visually plausible images, but tend to be unstable, difficult to control, and prone to hallucination of anatomy that does not correspond to any realistic background. Variational autoencoders (VAEs) introduced a probabilistic latent representation and more stable training [kingma2013auto, baur2021autoencoders, ehrhardt2022autoencoders], and structured or disentangled variants attempted to separate anatomy and pathology in latent space [chartsias2019disentangled, cetin2023attri]. However, because decoding still happens directly to full images, they often oversmooth fine pathological details and leak stochastic variation into non-lesional regions. More recently, diffusion models have become the dominant paradigm for high-fidelity synthesis in both natural and medical imaging [ho2020denoising, rombach2022high, pinaya2022brain, kazerouni2022diffusion, ahsan2025comprehensive]. By iteratively denoising from a simple prior, they capture rich, multimodal image statistics and offer strong sample diversity. However, most medical diffusion models still learn a distribution over entire images or over mask-conditioned images, without explicitly isolating where uncertainty should reside. Mask conditioning helps to specify where pathology should appear but not what it should look like internally [ren2019mask, zhang2024lefusion, meng2024multi]; lesions may have plausible boundaries, but lack realistic internal texture, heterogeneity, or progression patterns. As a result, even advanced models still entangle anatomical background with pathological variation and fail to exploit that anatomy is largely deterministic while pathology is a structured deviation (Fig. 1).
@@ 2 Related Work
Early GAN- and VAE-based models [goodfellow2020generative, isola2017image, wolterink2017deep, frid2018synthetic, yi2019generative, kingma2013auto, baur2021autoencoders, ehrhardt2022autoencoders] assume a holistic generative process $\mathbf{x}\sim p_{\theta}(\mathbf{x})$, which requires the network to assign the probability mass across the entire image manifold $\mathcal{X}\in\mathbb{R}^{H\times W\times C}$. This implicitly treats both the anatomical substrate $\mathbf{a}$ and the pathological appearance $\mathbf{d}$ as coupled random variables. Because $p_{\theta}(\mathbf{x})$ must account for the morphology of the lesion without explicit structural constraints, the model often exploits anatomical degrees of freedom to minimize the training objective, leading to realistic pathology at the cost of distorted background anatomy.
@@
To constrain the generative search space, mask-conditioned methods [ren2019mask, zhang2024lefusion, meng2024multi] narrow the problem to $\mathbf{x}\sim p_{\theta}(\mathbf{x}\mid\mathbf{m})$, where $\mathbf{m}\in\{0,1\}^{H\times W}$ defines a binary lesion region. Although this specifies where changes occur, it does not define how the disease should manifest. The mask $\mathbf{m}$ encodes spatial support but lacks internal semantic identity; consequently, the generator solves a boundary-respecting fill-in problem rather than a clinically grounded deviation. This produces plausible contours but insufficient internal heterogeneity, as $\mathbf{m}$ resolves spatial ambiguity but not distributional uncertainty of the state of the disease.
@@
Representation-based approaches introduce latent factors [bone2020learning], often decomposing the image into $(\mathbf{z}_{anat},\mathbf{z}_{path})$ to separate structural and pathological attributes. This aims to model $p_{\theta}(\mathbf{x})$ through a factorized prior $p(\mathbf{z}_{anat})p(\mathbf{z}_{path})$. However, the transformation back to the pixel space, $\mathbf{x}=f_{\theta}(\mathbf{z}_{anat},\mathbf{z}_{path})$, inevitably recouples these factors. In practice, stochastic variation intended for the pathology leaks into non-lesional regions. While the representation is disentangled in latent space, the generative mapping is not, meaning anatomical preservation is not guaranteed by construction.
@@ 3.3 Conditional Diffusion for Deviation Field Modeling
Building upon the decomposition defined in Eq. (1), we model the conditional distribution $p(\mathbf{r}\mid\mathbf{x}_{\mathrm{sub}},\mathbf{m})$ using a Denoising Diffusion Probabilistic Model (DDPM) [ho2020denoising]. The diffusion process is defined over the deviation field $\mathbf{r}$ rather than the global image manifold $\mathbf{x}$. This formulation ensures that the generative modeling capacity is exclusively allocated to the pathological appearance, while the anatomical substrate remains deterministic and invariant.
@@ 4.0.1 Dataset and Pre-processing
Experiments are conducted on the BraTS 2020 brain tumor dataset [menze2014multimodal, bakas2017advancing, bakas2018identifying], which comprises multi-institutional, multi-modal magnetic resonance images including T1, T1-weighted contrast-enhanced (T1Gd), T2-weighted and T2 Fluid Attenuated Inversion Recovery (FLAIR) volumes. The dataset provides expert voxel-level annotations for three tumor sub-regions: the necrotic and non-enhancing tumor core (NCR/NET), the peritumoral edema (ED), and the GD-enhancing tumor (ET). All imaging data are provided in a pre-processed format, including co-registration to a common anatomical template, interpolation to a uniform 1 $mm^{3}$ resolution, and skull-stripping.
@@
All scans undergo standardized preprocessing, including skull stripping, spatial resampling to a common grid, and normalization of the intensity. To disentangle the effect of the quality of the synthesis from that of the size of the training set, all the augmentation regimes, comprising the baseline methods (Brain-LDM [pinaya2022brain], MaskDiff-Inpaint [corneanu2024latentpaint]) are configured to contribute an equivalent number of synthetic samples to the training pool. To ensure a fair, controlled comparison and seamless integration with our proposed methods, we implement a pacth-wise VAE-GAN [gur2020hierarchical] within our unified PathoSyn framework. This allows us to standardize training protocols, architectural choices, and evaluation procedures across PathoSyn: VAE-GAN and our final model, PathoSyn: Diff, thus isolating the contribution of the generative paradigm rather than confounding factors from differing implementations.
@@ 4.2 Evaluation Framework
  • Segmentation: We benchmark across the supervised regimes (nnU-Net [isensee2021nnu], U-Net [ronneberger2015u]), semi-supervised regimes (Mean-Teacher U-Net [tarvainen2017mean]), and unsupervised (Contra-Seg [liu2023contrastive]). The results are reported as mean Dice similarity coefficient (DSC) $\pm$ standard deviation.
@@ 4.4 Mathematical Disentanglement Analysis
Structured Disentanglement via Deviation Space: Unlike dual-encoder frameworks (e.g., Meta-Auto [bone2020learning]) that still exhibit significant latent overlap, PathoSyn achieves near-orthogonality (C=0.072). By constraining the generation to a dedicated deviation space, we effectively “shield” the anatomical manifold from pathological stochasticity. This ensures that the digital twins generated are not only visually realistic, but also mathematically consistent, providing a safer and more reliable foundation for the training of clinical AI.
@@
Brain-LDM [pinaya2022brain] Holistic $\times$ 0.425 $\pm$ .082 1.24 $\pm$ .15
Meta-Auto [bone2020learning] Dual-Enc 0.245 $\pm$ .041 0.78 $\pm$ .12
@@ 4.5 Downstream Evaluation Protocol
Tab. 2 summarizes the segmentation performance (Dice Similarity Coefficient) across four distinct learning regimes, ranging from fully supervised to unsupervised configurations. The results indicate that the inclusion of PathoSyn: Diff synthetic data leads to a consistent and statistically significant improvement ($p<0.01$) in all evaluated architectures. Specifically, on the nnU-Net supervised baseline, PathoSyn: Diff achieves a mean Dice score of $0.845\pm 0.021$, outperforming the “No Augmentation” baseline by an absolute margin of $6.5\%$. As shown in Tab. 2, the performance gains are not limited to supervised models; in the unsupervised Contra-Seg regime, our diffusion-based approach yields a $5.5\%$ improvement over the “No Augmentation” baseline, suggesting that PathoSyn effectively captures the underlying pathological manifold in a way that remains useful even without explicit voxel-level labels. Furthermore, the reduced standard deviation observed in the PathoSyn cohorts suggests that our disentangled deviation-space modeling provides more consistent and stable training signals compared to holistic synthesis methods like Brain-LDM, which can occasionally introduce anatomical artifacts.
@@
Tab. 3 summarizes the impact of various synthesis strategies on the classification of pathological states in three diverse architectural backbones: ResNet-50 (CNN), DenseNet (3D-Volumetric) and Swin Transformer. The results indicate that data augmentation via PathoSyn: Diff consistently yields the highest discriminative performance, achieving a peak AUROC of 0.915 on the ResNet-50 backbone. This represents a substantial improvement over holistic synthesis models such as Brain-LDM and MaskDiff-Inpaint, underscoring the advantage of modeling stochasticity within a constrained deviation space rather than across the entire image manifold. PathoSyn: Diff markedly improves model calibration, as evidenced by the lowest observed Expected Calibration Error (ECE) across all test regimes (0.043–0.045). As indicated in Tab. 3, while conventional and holistic augmentations improve AUROC, they often provide marginal gains in ECE. In contrast, our diffusion-based approach reduces calibration error by approximately 35% compared to the No Augmentation baseline. This suggests that the high-fidelity, anatomically-grounded lesions generated by PathoSyn prevent the classifiers from becoming overconfident on synthetic artifacts, leading to more reliable and trustworthy probability estimates that are essential for clinical decision-support systems.
@@
Synthesis Method ResNet-50 (2D CNN) DenseNet (3D-Vol.) Swin Transformer
@@
Fig. 5 (Top Left) illustrates the tumor segmentation performance when models are augmented with various synthesis strategies. Beyond global improvements, we specifically analyze the sensitivity of segmentation models to the tumor scale. As shown in comparative analysis, while most generative methods provide marginal gains for large, high-contrast lesions, their performance typically degrades as the size of the lesion decreases due to the loss of high-frequency textural cues in the latent space. In contrast, PathoSyn: Diff demonstrates remarkable robustness across the entire size spectrum. Specifically, for small tumors (defined as the lower quartile of the volume distribution of the injury), our method achieves the most significant relative performance gain. This is attributed to our localized deviation-space modeling, which prevents the ”blurring” effect common in holistic models like Brain-LDM. By constraining the diffusion process to the specific pathological support m, PathoSyn preserves the sharp boundaries and mesoscopic textures necessary for a segmentation network to identify early-stage or small-scale pathologies. Tab. 2 and Fig. 5 collectively indicate that by providing high-fidelity synthetic examples of small lesions—which are often underrepresented in the training distribution-PathoSyn effectively mitigates the class imbalance at the voxel level, leading to a more calibrated and sensitive segmentation output.
@@ 4.6 Clinical Distributional Consistency
Fig. 4 qualitatively compares the clinical plausibility of the lesions generated. A primary requirement for clinical consistency in neuro-oncology is the adherence to anatomical boundaries and tissue-specific intensity profiles. While holistic models like Brain-LDM often generate “low-level” lesions that ignore the underlying textures, PathoSyn: Diff ensures that pathological deviations—such as peritumoral edema—respect the orientation of white matter tracts and the constraints of the ventricular system. Preservation of the anatomical substrate during the diffusion process ensures that the synthesized pathology appears physiologically ”anchored” rather than superimposed.
Fig. 5 (top right) quantifies the Deep Conceptual Texture Distance, serving as a computational proxy for high-level semantic alignment. In clinical practice, radiologists characterize gliomas through complex morphological patterns, including internal mottling, rim enhancement, and central necrosis. While holistic generative models often produce over-smoothed or biologically sterile textures, PathoSyn: Diff achieves the lowest conceptual distance to the authentic BRaTS distribution. This indicates that the model’s latent representations successfully capture the intricate semantic markers essential for diagnostic differentiation. The primary utility of this high-level alignment is the mitigation of shortcut learning. When synthetic data contains subtle digital artifacts or lacks biological complexity, downstream AI models may learn to recognize fake features rather than actual pathology. By narrowing the conceptual gap, PathoSyn ensures that a segmentation or classification network learns the same radiomic signatures present in human patients. This improves the clinical reliability and safety of AI tools, ensuring they generalize effectively to real-world diagnostic workflows without being misled by synthetic inconsistencies.
@@
Fig. 5 (bottom left) provides a granular evaluation of Intensity and Texture Consistency, encompassing first-order statistics (Mean, Skewness, Kurtosis) and second-order GLCM metrics (Contrast, Homogeneity). Radiologically, high-grade gliomas are defined by their extreme heterogeneity, characterized by high intensity variance and specific gradient magnitudes at the infiltrative margins. The alignment of our results with authentic clinical data across these descriptors confirms that PathoSyn effectively models the complex internal photometric environment of the lesion. This level of textural fidelity is essential to ensure that diagnostic models trained on these data are sensitive to the subtle variations of tissue used to determine tumor grade.
@@
Fig. 5 (bottom right) presents the Empirical Cumulative Distribution Function (CDF) of the global feature distance to the real pathological manifold. This curve provides a population-level assessment of generative reliability. The pronounced shift to the left of the PathoSyn curve indicates that the model consistently produces samples within the ”clinically acceptable” range of the authentic distribution. In a clinical context, this distributional stability confirms that the framework does not merely produce isolated high-quality samples but consistently populates the data space with diverse and biologically credible digital twins. This minimizes the risk of generating outlier samples that could degrade the robustness of downstream clinical AI applications.

### v2

@@ 1 Introduction
Pathological image generation is increasingly vital for medical image analysis, supporting data augmentation [11, 19, 22], disease progression modeling [39, 28, 29], and clinical decision support [27, 8]. However, acquiring balanced datasets is hindered by limited patient availability, ethical constraints, and high annotation costs. Generative models mitigate this scarcity by synthesizing realistic pathological images that complement real-world data. Beyond algorithmic utility, high-fidelity synthetic pathology aids in clinical intervention planning by illustrating plausible disease trajectories, surfacing rare edge cases, and enabling controlled evaluation of models prior to deployment [20, 33, 35, 39, 28]. Enhancing the fidelity and consistency of these images is therefore directly linked to developing safer, more reliable tools for clinical decision-making.
Despite growing interest, generating pathological images remains challenging due to the intrinsic asymmetry between anatomy and disease. The anatomical structure, including organ geometry and spatial layout, is largely stable for a given subject, whereas the pathological appearance is the main source of uncertainty, varying in intensity, texture, and clinical stage. The most clinically relevant variability is concentrated in the way disease perturbs a stable anatomical substrate. Existing generative models [22, 18] rarely make this distinction explicit. When operating directly in image space, anatomy and pathology are treated equally stochastic; the model must learn a high-dimensional distribution over all pixels simultaneously. This unnecessarily enlarges the solution space and permits the generator to introduce anatomical modifications in order to fit the data, thereby producing systematic distortions of anatomically stable structures. Consequently, current approaches fall into two suboptimal extremes: either they model the image as a holistic entity with no explicit separation [26], or they enforce complete separation via masking, which identifies location, but overlooks internal lesion appearance and continuity with surrounding tissue.
@@
Although deep learning-based methods have significantly advanced visual fidelity compared to traditional rule-based algorithms, they have remained largely confined to the same image-space paradigm [17, 37]. Early applications of generative adversarial networks (GANs) synthesized lesions by modifying entire slices or volumes, sometimes conditioned on labels or coarse masks [12, 16, 36, 11, 38]; these models can create sharp and visually plausible images, but tend to be unstable, difficult to control, and prone to hallucination of anatomy that does not correspond to any realistic background. Variational autoencoders (VAEs) introduced a probabilistic latent representation and more stable training [21, 4, 10], and structured or disentangled variants attempted to separate anatomy and pathology in latent space [7, 6]. However, because decoding still happens directly to full images, they often oversmooth fine pathological details and leak stochastic variation into non-lesional regions. More recently, diffusion models have become the dominant paradigm for high-fidelity synthesis in both natural and medical imaging [14, 31, 26, 18, 1]. By iteratively denoising from a simple prior, they capture rich, multimodal image statistics and offer strong sample diversity. However, most medical diffusion models still learn a distribution over entire images or over mask-conditioned images, without explicitly isolating where uncertainty should reside. Mask conditioning helps to specify where pathology should appear but not what it should look like internally [30, 40, 24]; lesions may have plausible boundaries, but lack realistic internal texture, heterogeneity, or progression patterns. As a result, even advanced models still entangle anatomical background with pathological variation and fail to exploit that anatomy is largely deterministic while pathology is a structured deviation (Fig. 1).
@@ 2 Related Work
Early GAN- and VAE-based models [12, 16, 36, 11, 38, 21, 4, 10] assume a holistic generative process $\mathbf{x}\sim p_{\theta}(\mathbf{x})$, which requires the network to assign the probability mass across the entire image manifold $\mathcal{X}\in\mathbb{R}^{H\times W\times C}$. This implicitly treats both the anatomical substrate $\mathbf{a}$ and the pathological appearance $\mathbf{d}$ as coupled random variables. Because $p_{\theta}(\mathbf{x})$ must account for the morphology of the lesion without explicit structural constraints, the model often exploits anatomical degrees of freedom to minimize the training objective, leading to realistic pathology at the cost of distorted background anatomy.
@@
To constrain the generative search space, mask-conditioned methods [30, 40, 24] narrow the problem to $\mathbf{x}\sim p_{\theta}(\mathbf{x}\mid\mathbf{m})$, where $\mathbf{m}\in\{0,1\}^{H\times W}$ defines a binary lesion region. Although this specifies where changes occur, it does not define how the disease should manifest. The mask $\mathbf{m}$ encodes spatial support but lacks internal semantic identity; consequently, the generator solves a boundary-respecting fill-in problem rather than a clinically grounded deviation. This produces plausible contours but insufficient internal heterogeneity, as $\mathbf{m}$ resolves spatial ambiguity but not distributional uncertainty of the state of the disease.
@@
Representation-based approaches introduce latent factors [5], often decomposing the image into $(\mathbf{z}_{anat},\mathbf{z}_{path})$ to separate structural and pathological attributes. This aims to model $p_{\theta}(\mathbf{x})$ through a factorized prior $p(\mathbf{z}_{anat})p(\mathbf{z}_{path})$. However, the transformation back to the pixel space, $\mathbf{x}=f_{\theta}(\mathbf{z}_{anat},\mathbf{z}_{path})$, inevitably recouples these factors. In practice, stochastic variation intended for the pathology leaks into non-lesional regions. While the representation is disentangled in latent space, the generative mapping is not, meaning anatomical preservation is not guaranteed by construction.
@@ 3.3 Conditional Diffusion for Deviation Field Modeling
Building upon the decomposition defined in Eq. (1), we model the conditional distribution $p(\mathbf{r}\mid\mathbf{x}_{\mathrm{sub}},\mathbf{m})$ using a Denoising Diffusion Probabilistic Model (DDPM) [14]. The diffusion process is defined over the deviation field $\mathbf{r}$ rather than the global image manifold $\mathbf{x}$. This formulation ensures that the generative modeling capacity is exclusively allocated to the pathological appearance, while the anatomical substrate remains deterministic and invariant.
@@ 4.0.1 Dataset and Pre-processing
Experiments are conducted on the BraTS 2020 brain tumor dataset [25, 2, 3], which comprises multi-institutional, multi-modal magnetic resonance images including T1, T1-weighted contrast-enhanced (T1Gd), T2-weighted and T2 Fluid Attenuated Inversion Recovery (FLAIR) volumes. The dataset provides expert voxel-level annotations for three tumor sub-regions: the necrotic and non-enhancing tumor core (NCR/NET), the peritumoral edema (ED), and the GD-enhancing tumor (ET). All imaging data are provided in a pre-processed format, including co-registration to a common anatomical template, interpolation to a uniform 1 $mm^{3}$ resolution, and skull-stripping.
@@
All scans undergo standardized preprocessing, including skull stripping, spatial resampling to a common grid, and normalization of the intensity. To disentangle the effect of the quality of the synthesis from that of the size of the training set, all the augmentation regimes, comprising the baseline methods (Brain-LDM [26], MaskDiff-Inpaint [9]) are configured to contribute an equivalent number of synthetic samples to the training pool. To ensure a fair, controlled comparison and seamless integration with our proposed methods, we implement a pacth-wise VAE-GAN [13] within our unified PathoSyn framework. This allows us to standardize training protocols, architectural choices, and evaluation procedures across PathoSyn: VAE-GAN and our final model, PathoSyn: Diff, thus isolating the contribution of the generative paradigm rather than confounding factors from differing implementations.
@@ 4.2 Evaluation Framework
  • Segmentation: We benchmark across the supervised regimes (nnU-Net [15], U-Net [32]), semi-supervised regimes (Mean-Teacher U-Net [34]), and unsupervised (Contra-Seg [23]). The results are reported as mean Dice similarity coefficient (DSC) $\pm$ standard deviation.
@@ 4.4 Mathematical Disentanglement Analysis
Structured Disentanglement via Deviation Space: Unlike dual-encoder frameworks (e.g., Meta-Auto [5]) that still exhibit significant latent overlap, PathoSyn achieves near-orthogonality (C=0.072). By constraining the generation to a dedicated deviation space, we effectively “shield” the anatomical manifold from pathological stochasticity. This ensures that the digital twins generated are not only visually realistic, but also mathematically consistent, providing a safer and more reliable foundation for the training of clinical AI.
@@
Brain-LDM [26] Holistic $\times$ 0.425 $\pm$ .082 1.24 $\pm$ .15
Meta-Auto [5] Dual-Enc 0.245 $\pm$ .041 0.78 $\pm$ .12
@@ 4.5 Downstream Evaluation Protocol
Tab. 2 summarizes Dice segmentation performance across four learning regimes, from fully supervised to unsupervised. Adding PathoSyn: Diff synthetic data yields consistent, statistically significant improvements ($p<0.01$) for all architectures. On the supervised nnU-Net baseline, PathoSyn: Diff reaches $0.845\pm 0.021$ Dice, an absolute gain of $6.5\%$ over “No Augmentation.” In the unsupervised Contra-Seg regime, it improves performance by $5.5\%$, indicating that PathoSyn effectively models the pathological manifold even without voxel-level labels. The lower standard deviation in PathoSyn cohorts further suggests that our disentangled deviation-space modeling provides more stable training signals than holistic methods like Brain-LDM, which may introduce anatomical artifacts.
@@
Tab. 3 summarizes the effect of different synthesis strategies on pathological state classification across three backbones: ResNet-50 (CNN), DenseNet (2D), and Swin Transformer. PathoSyn: Diff augmentation consistently yields the best performance, reaching a peak AUROC of 0.915 on ResNet-50, substantially outperforming holistic synthesis models such as Brain-LDM and MaskDiff-Inpaint. This highlights the benefit of modeling stochasticity within a constrained deviation space rather than across the full image manifold. PathoSyn: Diff also delivers the best calibration, with the lowest Expected Calibration Error (ECE) across all test regimes (0.043–0.045). While conventional and holistic augmentations increase AUROC, Tab. 3 shows they offer only modest ECE gains. In contrast, our diffusion-based approach reduces calibration error by about 35% relative to No Augmentation. This indicates that high-fidelity, anatomically grounded lesions from PathoSyn prevent overconfidence on synthetic artifacts, yielding more reliable probability estimates crucial for clinical decision support.
@@
Synthesis Method ResNet-50 (2D CNN) DenseNet (2D) Swin Transformer
@@
Fig. 5 (top left) shows tumor segmentation performance for models trained with different synthesis strategies. Beyond overall gains, we assess sensitivity to tumor size. Most generative methods offer modest improvements for large, high-contrast lesions but degrade sharply for smaller ones, largely due to loss of high-frequency texture in the latent space. In contrast, PathoSyn: Diff maintains strong performance across all lesion sizes. For small tumors (lowest quartile of lesion volume), it achieves the largest relative accuracy gain. This stems from its localized deviation-space modeling, which avoids the blurring seen in holistic models such as Brain-LDM. By restricting diffusion to the pathological support $m$, PathoSyn preserves sharp boundaries and mesoscopic textures essential for detecting early-stage or small lesions. Together, Tab. 2 and Fig. 5 show that by generating high-fidelity synthetic examples of typically underrepresented small lesions, PathoSyn mitigates voxel-level class imbalance and yields more calibrated, more sensitive tumor segmentation.
@@ 4.6 Clinical Distributional Consistency
Fig. 4 qualitatively compares the clinical plausibility of generated lesions. A key requirement in neuro-oncology is respecting anatomical boundaries and tissue-specific intensity profiles. Holistic models like Brain-LDM often generate “low-level” lesions that ignore underlying textures, whereas PathoSyn: Diff ensures that deviations—such as peritumoral edema—follow white matter tracts and ventricular constraints. Preserving the anatomical substrate during diffusion makes the synthesized pathology appear physiologically anchored rather than superimposed.
@@
Fig. 5 (top right) reports Intensity and Texture Consistency, including first-order statistics (Mean, Skewness, Kurtosis) and second-order GLCM metrics (Contrast, Homogeneity). High-grade gliomas are radiologically defined by pronounced heterogeneity, with high intensity variance and characteristic gradients at infiltrative margins. The close match between our results and clinical data across these descriptors shows that PathoSyn accurately models the lesion’s internal photometric structure. This textural fidelity is crucial for training diagnostic models that are sensitive to subtle tissue variations used to determine tumor grade.
Fig. 5 (bottom left) measures the Deep Conceptual Texture Distance as a proxy for high-level semantic alignment. Radiologists characterize gliomas by complex morphologies, including internal mottling, rim enhancement, and central necrosis. Holistic models tend to yield over-smoothed or biologically sterile textures, but PathoSyn: Diff achieves the smallest conceptual distance to the BRaTS distribution, indicating that its latent space captures the semantic markers critical for diagnosis. This high-level alignment helps prevent shortcut learning: if synthetic data contain digital artifacts or lack biological complexity, downstream models may learn fake cues instead of real pathology. By narrowing this conceptual gap, PathoSyn encourages segmentation and classification networks to learn the same radiomic signatures found in patients, improving clinical reliability and generalization to real-world workflows.
@@
Fig. 5 (bottom right) shows the Empirical CDF of the global feature distance to the real pathological manifold, providing a population-level view of generative reliability. The leftward shift of the PathoSyn curve indicates that samples consistently fall within the clinically acceptable region of the true distribution. Clinically, this distributional stability means the framework does not just produce occasional high-quality images but reliably generates diverse, biologically credible digital twins, reducing the risk of outliers that could undermine downstream clinical AI robustness.